# Exploring Representations and Interventions in Time Series Foundation Models

### v2

@@ Appendix B Additional Representation Similarity Metrics
To comprehensively analyze the similarity of learned representations, we considered several metrics commonly used in the literature. While our primary analysis relies on Centered Kernel Alignment (CKA) Kornblith et al. [2019], we also explored two additional similarity metrics, namely Cosine Similarity and Singular Vector Canonical Correlation Analysis (SVCCA) [Raghu et al., 2017]. Our findings consistently demonstrated similar patterns across all these metrics, underscoring the robustness of our results. We provide brief descriptions of Cosine Similarity and SVCCA below.
@@
Singular Vector Canonical Correlation Analysis (SVCCA) [Raghu et al., 2017]: SVCCA is a method that compares the similarity of representations by aligning subspaces spanned by the top singular vectors. It effectively reduces the dimensionality and then compares the correlations of the principal components. The SVCCA similarity between two activation matrices $\mathbf{X}$ and $\mathbf{Y}$ is computed as follows:
@@
Appendix C Model Descriptions and Specifications
C.1 MOMENT
@@ C.3 Moirai
Appendix D Additional Results
@@
(i) Periodicity, $\alpha=0$ (ii) Some trend, more periodicity, $\alpha=0.25$
@@
(iii) Trend and periodicity, $\alpha=0.5$ (iv) More trend, some periodicity, $\alpha=0.75$
(v) Trend, $\alpha=1.0$
Figure 13: Compositional Steering for Forecasting using the MOMENT model. The parameter $\alpha$ influences compositional steering, where both sinusoidal and increasing trend concepts are observed in the steered output when $\alpha$=0.5, showcasing how the steering technique can interpolate between different concept combinations in the latent space (ii). Steered output changes from sinusodial concepts to increasing trend concepts by varying $\alpha\in\{0.0,0.25,0.5,0.75,1.0\}$.
@@
Refer to caption Refer to caption
(i) Periodicity, $\alpha=0$ (ii) Some trend, more periodicity, $\alpha=0.25$
Refer to caption Refer to caption
(iii) Trend and periodicity, $\alpha=0.5$ (iv) More trend, some periodicity, $\alpha=0.75$
(v) Trend, $\alpha=1.0$
Figure 15: Compositional Steering for Forecasting using the Chronos model. The parameter $\alpha$ influences compositional steering, where both sinusoidal and increasing trend concepts are observed in the steered output when $\alpha$=0.5, showcasing how the steering technique can interpolate between different concept combinations in the latent space (ii). Steered output changes from sinusoidal concepts to increasing trend concepts by varying $\alpha\in\{0.0,0.25,0.5,0.75,1.0\}$.
D.1 Pruning experiments
@@
Table 1: Zero-shot imputation performance of MOMENT. Results averaged across four different masking rates: {12.5%, 25%, 37.5%, 50%} and five runs with different masking seeds. This table presents the Model Performance Metrics (Mean Absolute Error and Mean Squared Error) on a subset of the Time Series Pile [Goswami et al., 2024, Zhou et al., 2021]. The model names include: ”Vanilla” MOMENT-Large without any pruning, ”Block 1-3” for cases where only one block is pruned, and ”All” for all three blocks being pruned. The best results per dataset are bolded. For full results refer to 3
@@
Table 3: Zero-shot imputation performance of MOMENT. Results averaged across five runs with different masking seeds. This table presents the Model Performance Metrics (Mean Absolute Error and Mean Squared Error) on a subset of the Time Series Pile [Goswami et al., 2024, Zhou et al., 2021]. The model names include: ”Vanilla” MOMENT-Large without any pruning, ”Block 1-3” for cases where only one block is pruned, and ”All” for all three blocks being pruned.
Vanilla All Pruned
MAE MSE MAE MSE
Dataset Horizon
Exchange 96 0.236 0.109 0.240 0.113
192 0.333 0.215 0.335 0.218
336 0.474 0.417 0.460 0.394
720 0.757 1.003 0.780 1.066
ETTh1 96 0.409 0.385 0.409 0.388
192 0.426 0.411 0.426 0.414
336 0.437 0.423 0.437 0.424
720 0.464 0.443 0.470 0.460
ETTh2 96 0.346 0.287 0.351 0.296
192 0.386 0.350 0.389 0.356
336 0.409 0.370 0.413 0.382
720 0.440 0.404 0.439 0.404
ETTm1 96 0.347 0.290 0.346 0.29
192 0.373 0.330 0.369 0.326
336 0.386 0.352 0.386 0.354
720 0.420 0.409 0.422 0.414
ETTm2 96 0.260 0.171 0.262 0.173
192 0.300 0.231 0.304 0.236
336 0.337 0.287 0.343 0.294
720 0.392 0.372 0.395 0.372
ILI 24 1.307 3.260 1.219 2.981
36 1.347 3.516 1.286 3.209
48 1.405 3.828 1.374 3.479
60 1.428 3.989 1.387 3.602
Weather 96 0.209 0.153 0.209 0.152
192 0.248 0.197 0.248 0.198
336 0.285 0.246 0.287 0.247
720 0.337 0.316 0.337 0.317
Table 4: Fine-tuned forecasting performance of MOMENT. This table presents model performance metrics (Mean Absolute Error and Mean Squared Error) on a subset of the Time Series Pile [Goswami et al., 2024, Zhou et al., 2021]. Metrics are presented for MOMENT-Large without any pruning (Vanilla) and for all blocks pruned (All). Results are gathered from the best performance on the test set across 3 epochs of training with a batch size 64 and learning rate of 0.0001.

### v3

@@ Appendix B Additional Representation Similarity Metrics
To comprehensively analyze the similarity of learned representations, we considered several metrics commonly used in the literature. While our primary analysis relies on Centered Kernel Alignment (CKA) (Kornblith et al., 2019), we also explored two additional similarity metrics, namely Cosine Similarity and Singular Vector Canonical Correlation Analysis (SVCCA) (Raghu et al., 2017). Our findings consistently demonstrated similar patterns across all these metrics, underscoring the robustness of our results. We provide brief descriptions of Cosine Similarity and SVCCA below.
@@
Singular Vector Canonical Correlation Analysis (SVCCA) (Raghu et al., 2017): SVCCA is a method that compares the similarity of representations by aligning subspaces spanned by the top singular vectors. It effectively reduces the dimensionality and then compares the correlations of the principal components. The SVCCA similarity between two activation matrices $\mathbf{X}$ and $\mathbf{Y}$ is computed as follows:
@@
Appendix C Finding and Pruning Redundant Blocks in TSFMs
The term “block” refers to groups of consecutive layers within a transformer that exhibit high representational similarity. Consistent with prior work (Phang et al., 2021), we use “layer” to refer to individual transformer encoder or decoder blocks consistent with prior work, while ”block” refers to a higher-level structure made up of multiple such layers that share similar representations (Nguyen et al., 2021).
As an example, consider Figure 1 which illustrates the pairwise similarity between the 24 layers of MOMENT-Large. In this figure, lighter colors (yellow) represent higher representational similarity, as measured by Centered Kernel Alignment (CKA) (Kornblith et al., 2019). The identified blocks are outlined with red bounding boxes. For example, Block 1 comprises layers 1–5, Block 2 comprises layers 9–18, and Block 3 comprises layers 19–23.
Our pruning algorithm is summarized in Algorithm 1. Our pruning method involves retaining the first and last layers of each block while zeroing out the weights of the intermediate layers. This approach preserves the structural integrity of the block while leveraging the skip connections within transformer blocks to ensure that signals and gradients continue to flow through the network. In the case of MOMENT-Large’s Block 3, composed of layers 19–23, this means that layers 19 and 23 are retained, while the weights of layers 20, 21, and 22 are zeroed out. We have added a dedicated section on pruning in Appendix C to further clarify this process.
Algorithm 2 Block Identification and Filtering
0:  CKA similarity matrix $S$, similarity threshold $\tau$, minimum block size $k$
  Initialize $\text{blocks}\leftarrow\emptyset$
[…]
  Phase 1: Initial block identification
  for each encoder block $l_{i}$ do
     if $\forall l_{j}\in\text{current\_block}:\text{CKA}(l_{i},l_{j})\geq\tau$ then
[…]
     else
[…]
        end if
[…]
     end if
  end for
  Phase 2: Filter small blocks
[…]
  Phase 3: Verify block-wide self-similarity
[…]
     $\text{submatrix}\leftarrow S[\text{start}:\text{end},\text{start}:\text{end}]$
[…]
     end if
  end for
  return final_blocks
C.1 A Simple Algorithmic Approach to Identify Redundant Blocks
We identify redundant blocks in a TSFM through visual inspection, which aligns with prior work (Nguyen et al., 2021). Table 2 lists all the identified blocks in MOMENT-Large and Chronos-Large. In addition to visual inspection, redundant blocks can also be identified using algorithmic approaches.
Below we propose a simple algorithm to identify redundant blocks in TSFMs. First, we can define a block as:
$\text{Block}=\{l_{i},...,l_{j}\}$
where $\text{where }i<j\text{ and }l_{k}\text{ are adjacent transformer encoder blocks}$
Then, we can systematically identify these blocks using an algorithm that:
  1. 1.
    Identifies initial candidate blocks by grouping adjacent layers with pairwise CKA similarity above a threshold
  2. 2.
    Filters out blocks that are too small to be considered meaningful
  3. 3.
    Verifies that each block exhibits high similarity across all its constituent layers by examining the complete submatrix of similarities
C.2 Identified Blocks in MOMENT & Chronos
In this paper, identified redundant blocks visually. Below we specify, the redundant blocks for MOMENT-Large and Chronos-Large:
Models Block 1 Block 2 Block 3 Block 4
MOMENT-Large 1 – 5 9 – 18 19 –23 N/A
Chronos-Large 1 – 4 5 – 9 10 –13 15 – 22
Table 2: Visually identified blocks of redundant layers in MOMENT-Large and Chronos-Large.
Appendix D On Steering and the parameter $\lambda$
Below we provide some guidance on selecting good values of $\lambda$ and insights into its properties:
Selection and Impact
  • Optimal Range: Based on our empirical experiments, we found that the steering strength parameter $\lambda$ is most effective for interventions when its value lies within the interval $[0.1,2.0]$.
  • Lower Bound Considerations: Values of $\lambda<0.1$ often result in insufficient perturbation of the activation patterns, leading to suboptimal intervention effects that may not manifest visibly in the model’s output.
  • Upper Bound Effects: Setting $\lambda>2.0$ induces excessive perturbations that push activations beyond their typical distribution bounds, potentially resulting in degenerate or semantically meaningless outputs. In the PCA/latent space visualizations, these cases simply appear as more distant points along the steering direction.
Directional Properties
  • Reversibility: Multiplying the steering vector by -1 effectively reverses the direction of intervention, enabling bidirectional control (e.g., transforming concept $A\rightarrow B$ into $B\rightarrow A$).
  • Example application: For a steering vector trained to increase signal magnitude, applying its negative counterpart (-$\lambda\mathbf{S}$) produces controlled signal decrease, demonstrating the symmetric nature of the steering operation.
Practical Guidelines
  • Initial Calibration: We recommend starting with $\lambda=1.0$ and adjusting based on the observed intervention strength. In most cases value $\lambda=1.0$ works well and does not need tuning.
  • Task and Model-Specific Tuning: If $\lambda=1.0$ does not yield satisfactory results, the optimal value requires tuning based on both the specific steering objective and target model, necessitating empirical calibration to achieve the desired intervention strength.
  • Monitoring: When applying steering interventions, practitioners should monitor both the immediate output and latent space representations to ensure meaningful transformations while maintaining output coherence.
Appendix E Model Descriptions and Specifications
E.1 MOMENT
@@ E.3 Moirai
Appendix F Additional Results
F.1 Steering
@@
(i) Periodicity, $\beta=0$ (ii) Some trend, more periodicity, $\beta=0.25$
@@
(iii) Trend and periodicity, $\beta=0.5$ (iv) More trend, some periodicity, $\beta=0.75$
(v) Trend, $\beta=1.0$
Figure 12: Compositional Steering for Forecasting using the MOMENT model. The parameter $\beta$ influences compositional steering, where both sinusoidal and increasing trend concepts are observed in the steered output when $\beta$=0.5, showcasing how the steering technique can interpolate between different concept combinations in the latent space (ii). Steered output changes from sinusodial concepts to increasing trend concepts by varying $\beta\in\{0.0,0.25,0.5,0.75,1.0\}$.
Refer to caption Refer to caption
(i) Periodicity, $\beta=0$ (ii) Some trend, more periodicity, $\beta=0.25$
Refer to caption Refer to caption
(iii) Trend and periodicity, $\beta=0.5$ (iv) More trend, some periodicity, $\beta=0.75$
(v) Trend, $\beta=1.0$
Figure 13: Compositional Steering for Forecasting using the Chronos model. The parameter $\beta$ influences compositional steering, where both sinusoidal and increasing trend concepts are observed in the steered output when $\beta$=0.5, showcasing how the steering technique can interpolate between different concept combinations in the latent space (ii). Steered output changes from sinusoidal concepts to increasing trend concepts by varying $\beta\in\{0.0,0.25,0.5,0.75,1.0\}$.
Application of Concept Steering on a Real-World Dataset
We demonstrate the practical utility of concept steering on the ECG5000 dataset, which contains electrocardiogram readings classified as either normal or abnormal heart patterns. Using a MOMENT with an SVM classifier, we achieve strong baseline performance in distinguishing between the two classes.
For the steering experiment, we compute steering matrices using the median method to capture the concept difference between normal and abnormal patterns. Analysis in the activation space reveals that our steering approach successfully moves samples toward the target concept (Figure 14). This bidirectional movement is evident the PCA visualization of activation patterns.
To validate our approach quantitatively, we applied the steering transformations to 30 samples. While these samples were initially classified correctly with 100% accuracy as normal heartbeats, after steering, all of these samples swapped output classes, confirming successful concept transfer. These results suggest that concept steering can effectively capture and manipulate clinically relevant patterns in physiological data.
PCA Visualizations
Refer to caption Refer to caption
(a) Normal to Abnormal (b) Abnormal to Normal
Figure 14: Visualization of Concept Steering on ECG Data. The figure shows PCA projections of the activation space, where arrows indicate the direction of steering. Blue and red points represent normal (class 0) and abnormal (class 1) samples respectively, with green points showing the original samples and orange points showing their steered versions.
F.2 Pruning
@@
Chronos Family
(i) Tiny (ii) Mini (iii) Small (iv) Base (v) Large
Figure 16: How does model size influence the patterns of learned representations? Smaller models exhibit more discrete and distinct block structures, while larger models (Base and Large) display increasingly intricate and clustered patterns, reflecting more nuanced and gradual transformations across layers. Notably, the emergence of blocks-like patterns in the Large model appears unpredictable from patterns observed in smaller models.
@@
Table 3: Zero-shot imputation performance of MOMENT. Results averaged across four different masking rates: {12.5%, 25%, 37.5%, 50%} and five runs with different masking seeds. This table presents the Model Performance Metrics (Mean Absolute Error and Mean Squared Error) on a subset of the Time Series Pile (Goswami et al., 2024; Zhou et al., 2021). The model names include: ”Vanilla” MOMENT-Large without any pruning, ”Block 1-3” for cases where only one block is pruned, and ”All” for all three blocks being pruned. The best results per dataset are bolded. For full results refer to 5
@@
Table 5: Zero-shot imputation performance of MOMENT. Results averaged across five runs with different masking seeds. This table presents the Model Performance Metrics (Mean Absolute Error and Mean Squared Error) on a subset of the Time Series Pile (Goswami et al., 2024; Zhou et al., 2021). The model names include: ”Vanilla” MOMENT-Large without any pruning, ”Block 1-3” for cases where only one block is pruned, and ”All” for all three blocks being pruned.
Horizon Horizon
Dataset 96 192 336 720 96 192 336 720
Vanilla MAE All Pruned MAE
ETTh1 0.409 0.426 0.437 0.464 0.409 0.426 0.437 0.470
ETTh2 0.346 0.386 0.409 0.440 0.351 0.389 0.413 0.439
ETTm1 0.347 0.373 0.386 0.420 0.346 0.369 0.386 0.422
ETTm2 0.260 0.300 0.337 0.392 0.262 0.304 0.343 0.395
Weather 0.209 0.248 0.285 0.337 0.209 0.248 0.287 0.337
Vanilla MSE All Pruned MSE
ETTh1 0.385 0.411 0.423 0.443 0.388 0.414 0.424 0.460
ETTh2 0.287 0.350 0.370 0.404 0.296 0.356 0.382 0.404
ETTm1 0.290 0.330 0.352 0.409 0.29 0.326 0.354 0.414
ETTm2 0.171 0.231 0.287 0.372 0.173 0.236 0.294 0.372
Weather 0.153 0.197 0.246 0.316 0.152 0.198 0.247 0.317
Table 6: Fine-tuned forecasting performance of MOMENT. This table presents model performance metrics (Mean Absolute Error and Mean Squared Error) on a subset of the Time Series Pile (Goswami et al., 2024; Zhou et al., 2021). Metrics are presented for MOMENT-Large without any pruning (Vanilla) and for all blocks pruned (All). Results are gathered from the best performance on the test set across 3 epochs of training with a batch size 64 and learning rate of 0.0001.